OpenStudy (anonymous):

prove by induction that ((x^(k))-1)=(x-1)(x^(k-1)+x^(k-2)+...+x+1) i got stuck when i have to prove k=n+1 from k=n being assumed true

OpenStudy (anonymous):

something is wrong with that formula

OpenStudy (anonymous):

what is wrong with the formula?

OpenStudy (kinggeorge):

Looks correct to me.

OpenStudy (kinggeorge):

Anyways, you have to start with a base case of k=1. Then, you have\[(x-1)\overset{\text{?}}{=} (x-1)(1)\]Which is true, so the base case is good.

OpenStudy (anonymous):

yup i have completed the base case and i got stuck when i assume k=n is true and i want to prove k=n+1 from k=n

OpenStudy (kinggeorge):

Now you assume true up to some \(n\in\mathbb{Z}^+\). So \[(x^n-1)=(x-1)(x^{n-1}+x^{n-2}+...+x+1)\]

OpenStudy (kinggeorge):

Now, look at \[(x-1)(x^n+x^{n-1}+x^{n-2}+...+x+1).\]This is equal to \[(x-1)x^n+(x-1)(x^{n-1}+x^{n-2}+...+x+1)\]which by our inductive hypothesis, is equal to \[(x-1)x^n+(x^n-1)=x^{n+1}-x^n+x^n-1=x^{n+1}-1.\]

OpenStudy (kinggeorge):

That make sense?

OpenStudy (anonymous):

ok that makes a lot of sense. thank you so much!

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

oh i read the exponent wrong, sorry